OpenStudy (anonymous):

Please Help ASAP !! How can the properties of exponents help solve exponential equations?

OpenStudy (anonymous):

the properties of exponents, as in logs? lf you have: \[\ln(5^3)=3\ln(5)\]

OpenStudy (anonymous):

Thats all it says its a very confusing questions i have more if you would be willing to help they are worded badly i need help clarifying them

OpenStudy (anonymous):

Well I guess the short answer would be: Exponential Functions can be solved more easily by using logarithms, as defined by the properties of all exponents.

OpenStudy (anonymous):

I think thats the correts answer these questions hate me but i appreciate your help could you help me with one more ?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

just a moment

OpenStudy (anonymous):

How can the properties of exponents help solve logarithmic equations?

OpenStudy (anonymous):

@nfgallimore

OpenStudy (anonymous):

I would say logarithmic equations are exponential functions

OpenStudy (anonymous):

Same question basically

OpenStudy (anonymous):

Do you know how logarithmic functions are created?

OpenStudy (anonymous):

I thought exponential was the inverse of logarithmic am i wrong ?

OpenStudy (anonymous):

@nfgallimore